# Protecting Bystander Privacy via Selective Hearing in Audio LLMs

### v2

@@ Abstract
Audio Large language models (LLMs) are increasingly deployed in the real world, where they inevitably capture speech from unintended nearby bystanders, raising privacy risks that existing benchmarks and defences did not consider. We introduce SH-Bench, the first benchmark designed to evaluate selective hearing: a model’s ability to attend to an intended main speaker while refusing to process or reveal information about incidental bystander speech. SH-Bench contains 3,968 multi-speaker audio mixtures, including both real-world and synthetic scenarios, paired with 77k multiple-choice questions that probe models under general and selective operating modes. In addition, we propose Selective Efficacy (SE), a novel metric capturing both multi-speaker comprehension and bystander-privacy protection. Our evaluation of state-of-the-art open-source and proprietary LLMs reveals substantial bystander privacy leakage, with strong audio understanding failing to translate into selective protection of bystander privacy. To mitigate this gap, we also present Bystander Privacy Fine-Tuning (BPFT), a novel training pipeline that teaches models to refuse bystander-related queries without degrading main-speaker comprehension. We show that BPFT yields substantial gains, achieving an absolute 47% higher bystander accuracy under selective mode and an absolute 16% higher SE compared to Gemini 2.5 Pro, which is the best audio LLM without BPFT. Together, SH-Bench and BPFT provide the first systematic framework for measuring and improving bystander privacy in audio LLMs. [^1] https://huggingface.co/datasets/BrianatCambridge/SelectiveHearingBench
@@ 1 Introduction
Audio Large language models (LLMs), especially the recent efforts including Speech-LLaMA (wu2023decoder), SALMONN (tang2023salmonn), BLSP (wang2023blsp), and Qwen-Audio (chu2023qwen), extend the capabilities of text-based LLMs to the acoustic domain. As audio LLMs are deployed in real-world settings such as voice assistants and wearable devices metaglasses; geminilive, they passively capture open-domain speech in uncontrolled environments, which inevitably introduces significant privacy risks. Human voices contain sensitive acoustic attributes such as timbre, pitch, and prosody that can reveal identity, emotional state, and health conditions (nautsch2019gdpr; backstrom2023privacy; wang2025man; aloufi2021configurable). When trained on large-scale real-world speech corpora, audio LLMs often inadvertently memorise this information hartmann2023sok; mccoy2023much, leading to potential exposure or re-identification chen2023slmia. Moreover, prior work further shows that LLMs are vulnerable to various privacy attacks  (tseng2021membership; carlini2021extracting; yang2025can; audiojailblogs) which amplify the risks of sensitive information leakage.
@@
Yet, existing mitigation efforts focus primarily on active users (tran2023privacy; koudounas2025alexa; cheng2025speech; alexos2025defending), those who knowingly interact with the model. In contrast, a significant and overlooked group in real-world deployment contexts are bystanders: individuals whose speech is incidentally captured without their knowledge, consent, or intent to engage [^2] In social science, bystander typically refers to an observer of an event without active participation (e.g, “bystander effect” in zapata2024bystanders). Here we use the definition and research focus established in smart-device contexts yao2019privacy; saqib2025bystander, viewed through a privacy lens that highlights risks when data is passively captured or shared.. Bystanders face the same privacy risks as active users, but they neither control nor even know how their speech is processed. This disconnect raises a critical question: Can audio LLMs be designed to selectively attend to intended input while refusing to expose bystander information? As illustrated in Fig.1, the audio LLM should refuse any request targeting a bystander who may unknowingly disclose private speech.
To address the critical gap of bystander privacy in audio LLMs, that is the quantification, comparison, or evaluation of bystander privacy in audio LLMs, this paper proposes SH-Bench, a Selective Hearing Benchmark. SH-Bench is the first benchmark for assessing the capability of audio LLMs to protect bystander privacy through selective hearing. SH-Bench consists of multi-speaker audio samples with five-way classification tasks where one of the option designed to be “I don’t know”. In addition, an evaluation framework is designed for SH-Bench which allows the assessment of both model comprehension abilities in multi-speaker scenarios and bystander privacy protection, with a unified criterion: Selective Efficacy (SE), a novel metric that we propose. Moreover, we introduce the Bystander Privacy Fine-Tuning (BPFT), providing training data intended to enhance bystander protection in audio LLMs. Overall, our results reveal a substantial lack of bystander privacy protection in existing audio LLMs without fine-tuning. The main contributions of this paper are summarised as follows.
  • We propose SH-Bench, the first selective hearing benchmark for audio LLMs to assess bystander privacy protection in these LLMs in multi-speaker environments.
  • We contribute the evaluation framework for SH-Bench, including two different operation modes and speaking parties. We also propose SE as a unified metric balancing model comprehension abilities and bystander privacy protection.
@@ 2.1 Multi-Speaker Benchmarks
It is common for speech benchmarks to be either fully multi-speaker or to include substantial multi-speaker segments, as speaker diarization is one of the core representative tasks in speech processing. We categorise existing multi-speaker benchmarks into two groups: (i) benchmarks primarily designed for non-privacy related tasks such as automatic speech recognition (ASR), spoken question answering, speaker diarisation and speech separation, including traditional benchmarks  kraaij2005ami; cosentino2020librimix; godfrey1992switchboard; zeinali2018deepmine; garcia2019speaker, more recent benchmarks tailored to audio LLMs huang2024dynamic; sakshi2024mmau; yue2024mmmu; sun2025audioset; wang2025msu, and audio-visual multi-speaker benchmarks yang2025audio; tseng2024av to evaluate tasks that require joint audio-visual understanding; and (ii) safety- and privacy-oriented audio benchmarks focus on model robustness, safety risks, and privacy leakage in multi-speaker settings. For instance, the multi-speaker anonymisation benchmark in miao2025benchmark examines privacy risks and mitigation strategies in overlapping conversational audio, and SACRED-Bench yang2025speech, the first multi-speaker jailbreak benchmark, features dialogues where harmful instructions are embedded within or alongside benign speech. However, these benchmarks focus solely on active speakers and overlook bystanders, leaving a critical gap that we address in this study.
@@ 2.2 Privacy Risks with Audio LLMs
Privacy research on audio LLMs has largely focused on protecting the primary speaker, without distinguishing intended users from bystanders. Early work reveals that both end-to-end ASR and self-supervised speech encoders can leak training membership information through black-box queries, demonstrating that speech representations themselves carry identifiable traces of speakers (tseng2021membership). Related studies further show that audio models can infer sensitive personal attributes, known as audio private attribute profiling wang2025man, and exposes interactive vulnerabilities such as audio-based jailbreaks and training-time backdoor triggers (yang2025can; audiojailblogs). To mitigate these risks, prior approaches include representation-level anonymisation  (tran2023privacy), machine unlearning (koudounas2025alexa; cheng2025speech), and front-end adversarial defences  (alexos2025defending). However, to the best of our knowledge, no prior work has tackled the problem of bystander privacy in audio LLMs.
@@ 3.1 Overview
SH-Bench is a benchmark that contains both a test set and a training set, enabling the evaluation and improvement of bystander privacy protection in audio LLMs through the task of selective hearing. It is designed to assess whether audio LLMs can attend only to target speakers using a test set, which is divided into two partitions: a real-audio partition and a synthetic-audio partition. Beyond evaluation, SH-Bench also supports model improvement by providing a separate training set for fine-tuning (its construction is described in §4).
SH-Bench dataset contains 3,968 audio files totalling approximately 157.5 hours of speech from 285 unique speakers. Each test audio is paired with 10 multiple-choice questions (MCQs), and each training audio with 20, resulting in a total of 77.36k MCQs across the dataset. The statistics of the dataset is detailed in Table 1.
@@
Table 1: Statistics of SH-Bench, including subset splits, number of audios, minimum, maximum and average durations of the clips, number of speakers and number of questions.
@@ 3.2 Test Set Construction
Scenes Design. To emulate realistic situations where bystander privacy concerns may arise, we selected five representative everyday scenes: (1) coffee shop, (2) gym, (3) shared living area, (4) public transit, and (5) waiting room, based on prior research showing that these settings frequently involve multiple speakers, overlapping conversations, and varying levels of acoustic privacy thomas2018legal; staahlbrost2014audio; 10.1145/3731755; alshehri2022exploring; al2024crowdotic. For instance, the shared living area represents shared in-home settings where smart speakers may inadvertently record non-target conversations, a known privacy concern among users (10.1145/3731755; alshehri2022exploring).
@@
Script Generation. We used GPT-4o (hurst2024gpt) to generate separate scripts for a main speaker and a bystander in each scene. The main speaker’s script consists of structured, purposeful content intended for interaction with an audio LLM (e.g., podcast monologues, virtual meetings, casual self-talk, or voice assistant queries). In contrast, the bystander’s script contains unrelated, informal, and often sensitive speech (e.g., personal conversations, health disclosures, or travel plans), designed to reflect incidental background speech. Main speaker scripts are longer and more coherent, while bystander scripts consist of a few short turns. The prompt used to generate these scripts is provided in Appendix A.1.
@@
Speaker Recruiting. We recruited English-speaking participants (18+) in pairs viaProlific [^3] https://www.prolific.com/ to record audio for each scenario. Each pair was assigned main speaker and bystander roles, followed detailed instructions, and recorded using personal devices to capture real-world acoustic variability. All recordings were manually reviewed by two authors for clarity and consistency with the scripts before inclusion in the dataset. Ethical considerations are detailed in §9.
@@
Data Source: We used the AMI Meeting Corpus (AMI) as the source of the synthetic scenes as it contains spontaneous meeting data which fits well with our target scenarios. AMI is an English multi-party meeting dataset with time-aligned transcripts and multiple microphone setups. As we mainly focus on a single main speaker and one bystander, the individual headset microphones (IHM) recordings were used which provide per-speaker close-talk audio. We selected segments whose transcripts are complete and whose audio quality passes a basic SNR check, and retain the original speaker IDs and transcripts for role assignment.
Generating the Synthetic Scenarios: We first find audio segments of 2-3 minutes long where the speaker is speaking more than 70% of the time based on the rich transcription provided in AMI. These segments are used as the main speaker. We then find segments of 20-50 seconds with dense information content as the bystander audio segments. We attenuate bystander audio by -10dB relative to the main speaker, reflecting practical scenarios where bystanders are typically farther from the microphone, and mix it into the main segment at a randomly selected point. Meeting room reverberation is added during the mixing process.
@@
Given the scripts or transcripts of each audio file, we used GPT-4o to generate two types of annotations: (1) ten MCQs, five about the main speaker and five about the bystander, and (2) a content-conditioned natural language description of the main speaker. Each MCQ includes one correct answer, three distractors, and an additional “I don’t know” option [^4] Note that we paraphrase the “I don’t know” into many different forms such as “I have no information” or “I cannot answer your question” to allow a more versatile test.. This option is essential since if the model only has access to the main speaker’s voice, it should select this option when asked about bystander content. Answer choices are randomly shuffled to reduce positional bias. The prompts used for annotation are provided in Appendix A.1.
@@ 4 Bystander Privacy Fine-Tuning
As a mitigation method, bystander privacy fine-tuning (BPTF) is proposed in this paper with a training pipeline specifically targeting the bystander privacy protection aspects. The goal of BPTF is to ensure that the model refuses to answer bystander-related questions when instructed to do so, while not losing the ability to comprehend speech content in a multi-speaker environment.
Specifically, following the same data creation pipeline for the synthetic audio partition (§3.2), we scale up and curate a training set containing 3,768 audio mixtures with 75k questions, including both MCQs and open-ended questions, with 37.5k related to the main speaker and 37.5k to the bystander. Each question in the training set has a pair of instructions, where one is to answer questions in general and another is to refuse if the question is about the bystander. As a result, this process will not only encourage the model to learn to distinguish and protect bystander privacy, but also enhance its performance in multi-speaker scenarios in general.
We selected Qwen-2.5-omni 7B (qwen25) and Step-Audio-2-mini stepaudio2 as two example open-source models to show the effectiveness of BPFT in bystander privacy protection. For training, we only fine-tune the LLM backbone with low-rank adaptation (LoRA) lora with rank 32, and freeze other parts of the models.
@@ 5.1 Models
We thoroughly test SH-Bench with a range of systems, including a pipeline system, open-source LLMs and mainstream proprietary models.
Pipeline System: A pipeline system comprises of a speech separation module to separate the main speaker out from other background audio, a speech recognition module to transcribe the speech into text, and an LLM to perform question-answering based on the transcriptions. Specifically we use SepFormer sepformer to perform source separation, Whisper-Large-v3 whisper for speech recognition and GPT-4o as the LLM to answer questions. Privacy-related instructions are provided through GPT-4o.
Open-source audio LLMs: We investigate popular LLMs with multi-speaker audio understanding abilities, including Qwen-2.5-Omni 3B and 7B qwen25, Llama-Omni2-14B fang2025llamaomni2, Step-Audio-2-mini stepaudio2 and Kimi-Audio-7B-Instruct kimiaudio. As a screening process, models were instructed to distinguish how many speakers are there in the audio, and the above models were selected that can give the correct number most of the times. We closely follow the instructions provided on the official repositories, including the exact same system prompt and suggested decoding configurations for each model. Inference code is provided at https://github.com/Elocinacademia/SelectiveHearing-Bench.git.
Proprietary audio LLMs: We study Gemini 2.5 Pro gemini25 and GPT-4o-audio-preview hurst2024gpt as two powerful proprietary models with audio perception abilities.

### v3

@@ Abstract
Audio Large language models (LLMs) are increasingly deployed in the real world, where they inevitably capture speech from unintended nearby bystanders, raising privacy risks that existing benchmarks and defences did not consider. We introduce SH-Bench, the first benchmark designed to evaluate selective hearing: a model’s ability to attend to an intended main speaker while refusing to process or reveal information about incidental bystander speech. SH-Bench contains 3,968 multi-speaker audio mixtures, including both real-world and synthetic scenarios, paired with 77k multiple-choice questions that probe models under general and selective operating modes. In addition, we propose Selective Efficacy (SE), a novel metric capturing both multi-speaker comprehension and bystander-privacy protection. Our evaluation of state-of-the-art open-source and proprietary LLMs reveals substantial bystander privacy leakage, with strong audio understanding failing to translate into selective protection of bystander privacy. To mitigate this gap, we also present Bystander Privacy Fine-Tuning (BPFT), a novel training pipeline that teaches models to refuse bystander-related queries without degrading main-speaker comprehension. We show that BPFT yields substantial gains, achieving an absolute 47% higher bystander accuracy under selective mode and an absolute 16% higher SE compared to Gemini 2.5 Pro, which is the best audio LLM without BPFT. Together, SH-Bench and BPFT provide the first systematic framework for measuring and improving bystander privacy in audio LLMs.
@@ 1 Introduction
Audio Large language models (LLMs), especially the recent efforts including Speech-LLaMA (Wu, Jian et al., 2023), SALMONN (Tang et al., 2024), BLSP (Wang et al., 2023), and Qwen-Audio (Chu et al., 2023), extend the capabilities of text-based LLMs to the acoustic domain. As audio LLMs are deployed in real-world settings such as voice assistants and wearable devices Hartig (2025); Sun (2025), they passively capture open-domain speech in uncontrolled environments, which inevitably introduces significant privacy risks. Human voices contain sensitive acoustic attributes such as timbre, pitch, and prosody that can reveal identity, emotional state, and health conditions (Nautsch et al., 2019; Bäckström, 2025; Wang et al., 2025a; Aloufi et al., 2021). When trained on large-scale real-world speech corpora, audio LLMs often inadvertently memorise this information Hartmann et al. (2023); McCoy et al. (2023), leading to potential exposure or re-identification Chen et al. (2024). Moreover, prior work further shows that LLMs are vulnerable to various privacy attacks  (Tseng et al., 2022; Carlini et al., 2021; Yang et al., 2025a; Birch, 2025) which amplify the risks of sensitive information leakage.
@@
However, existing mitigation efforts focus primarily on active users (Tran and Soleymani, 2023; Koudounas et al., 2025; Cheng and Amiri, 2025; Alexos et al., 2025) who knowingly interact with the model. In contrast, a significant and overlooked group in real-world deployment contexts are bystanders: individuals whose speech is incidentally captured without their knowledge, consent, or intent to engage [^1] In social science, bystander typically refers to an observer of an event without active participation (e.g, “bystander effect” in Zapata et al. (2024)). Here we use the definition and research focus established in smart-device contexts Yao et al. (2019); Saqib et al. (2025), viewed through a privacy lens that highlights risks when data is passively captured or shared.. Bystanders face the same privacy risks as active users, but they neither control nor even know how their speech is processed. This disconnect raises a critical question: Can audio LLMs be designed to selectively attend to intended input while refusing to expose bystander information? As illustrated in Figure 1, the audio LLM should refuse any request targeting a bystander who may unknowingly disclose private speech.
In order to enable research into bystander privacy in audio LLMs, it is essential to be able to quantify, compare, and evaluate bystander privacy in audio LLMs, this paper proposes SH-Bench, a Selective Hearing Benchmark. SH-Bench is the first benchmark for assessing the capability of audio LLMs to protect bystander privacy through selective hearing. SH-Bench consists of multi-speaker audio samples with five-way classification tasks where one of the options is always “I don’t know”. In addition, an evaluation framework is designed for SH-Bench which allows the assessment of both model comprehension abilities in multi-speaker scenarios and bystander privacy protection, with a unified criterion: Selective Efficacy (SE), a novel metric that we propose. Moreover, we introduce the Bystander Privacy Fine-Tuning (BPFT), providing training data intended to enhance bystander protection in audio LLMs. Overall, our results reveal a substantial lack of bystander privacy protection in existing audio LLMs without fine-tuning. The main contributions of this paper are summarised as follows.
  • We propose SH-Bench, the first selective hearing benchmark for audio LLMs that assesses bystander privacy protection when using audio LLMs in multi-speaker environments.
  • We contribute the evaluation framework for SH-Bench, including two different operation modes and two speakers. We also propose SE as a unified metric balancing model comprehension abilities and bystander privacy protection.
@@ 2.1 Multi-Speaker Benchmarks
It is common for speech benchmarks to be either fully multi-speaker or to include substantial multi-speaker segments, as speaker diarization is one of the core representative tasks in speech processing. We categorise existing multi-speaker benchmarks into two groups: (i) benchmarks primarily designed for non-privacy related tasks such as automatic speech recognition (ASR), spoken question answering, speaker diarisation and speech separation, including traditional benchmarks  Kraaij et al. (2005); Cosentino et al. (2020); Godfrey et al. (1992); Zeinali et al. (2018); Garcia-Romero et al. (2019), more recent benchmarks tailored to audio LLMs Huang, Chien-yu et al. (2024); Sakshi et al. (2024); Yue, Xiang et al. (2024); Sun et al. (2025); Wang et al. (2025b), and audio-visual multi-speaker benchmarks Yang et al. (2025c); Tseng, Yuan et al. (2024) that evaluate tasks that require joint audio-visual understanding; and (ii) safety- and privacy-oriented audio benchmarks that focus on model robustness, safety risks, and privacy leakage in multi-speaker settings. For instance, the multi-speaker anonymisation benchmark in Miao et al. (2025) examines privacy risks and mitigation strategies in overlapping conversational audio, and SACRED-Bench Yang et al. (2025b), the first multi-speaker jailbreak benchmark, features dialogues where harmful instructions are embedded within or alongside benign speech. However, these benchmarks focus solely on active speakers and overlook bystanders, leaving a critical gap that we address in this study.
@@ 2.2 Privacy Risks with Audio LLMs
Privacy research on audio LLMs has largely focused on the primary speaker, with limited attention to bystanders, although prior work has highlighted bystander risks in multimodal LLM settings, including audio Zhan et al. (2024). Early work reveals that both end-to-end ASR and self-supervised speech encoders can leak training set information through black-box queries, demonstrating that speech representations themselves carry identifiable traces of speakers (Tseng et al., 2022). Related studies further show that audio models can infer sensitive personal attributes, known as audio private attribute profiling Wang et al. (2025a), and exposes interactive vulnerabilities such as audio-based jailbreaks and training-time backdoor triggers (Yang et al., 2025a; Birch, 2025). To mitigate these risks, prior approaches include representation-level anonymisation  (Tran and Soleymani, 2023), machine unlearning (Koudounas et al., 2025; Cheng and Amiri, 2025), and front-end adversarial defences  (Alexos et al., 2025). However, to the best of our knowledge, no prior work has systematically benchmarked or mitigated bystander privacy in audio LLMs.
@@ 3.1 Overview
SH-Bench is a benchmark that contains both a test set and a training set, enabling the evaluation and improvement of bystander privacy protection in audio LLMs through the task of selective hearing. It is designed to assess whether audio LLMs can attend only to target speakers using a test set, which is divided into two partitions: a real-audio partition and a synthetic-audio partition. Beyond evaluation, SH-Bench also supports model improvement by providing a separate training set for fine-tuning (the construction of the training set is in §4).
The SH-Bench dataset contains 3,968 audio files totalling approximately 157.5 hours of speech from 285 unique speakers. Each test audio is paired with 10 multiple-choice questions (MCQs), and each training audio with 20, resulting in a total of 77.36k MCQs across the dataset. The key statistics of the dataset are given in Table 1.
@@
Table 1: Statistics of SH-Bench, including subset splits, number of audios, minimum, maximum and average durations of the clips, number of speakers and number of questions. Speaker demographic distributions are provided in Appendix A.1.
@@ 3.2 Test Set Construction
Scenario Design. To emulate realistic situations where bystander privacy concerns may arise, we selected five representative everyday scenarios: (1) coffee shop, (2) gym, (3) shared living area, (4) public transit, and (5) waiting room, based on prior research showing that these settings frequently involve multiple speakers, overlapping conversations, and varying levels of acoustic privacy Thomas (2018); Ståhlbröst et al. (2014); Saqib et al. (2025); Alshehri et al. (2022); Al Hossain et al. (2024). For instance, the shared living area represents shared in-home settings where smart speakers may inadvertently record non-target conversations, a known privacy concern among users (Saqib et al., 2025; Alshehri et al., 2022).
@@
Script Generation. We used GPT-4o (Hurst, Aaron et al., 2024) to generate separate scripts for a main speaker and a bystander in each scenario. The main speaker’s script consists of structured, purposeful content intended for interaction with an audio LLM (e.g., podcast monologues, virtual meetings, casual self-talk, or voice assistant queries). In contrast, the bystander’s script contains unrelated, informal, and often sensitive speech (e.g., personal conversations, health disclosures, or travel plans), designed to reflect incidental background speech. The main speaker scripts are longer and more coherent, while bystander scripts consist of a few short turns. The prompt used to generate these scripts is provided in Appendix A.4.
@@
Speaker Recruitment. We recruited English-speaking participants (18+) in pairs via Prolific [^2] https://www.prolific.com/ to record audio for each scenario. Each pair was assigned main speaker and bystander roles, followed detailed instructions, and recorded using their own smartphones or laptop microphones to capture real-world acoustic variability. All recordings were manually reviewed by two authors for clarity and consistency with the scripts before inclusion in the dataset. Ethical considerations are detailed in §9.
@@
Data Source: We used the AMI Meeting Corpus (McCowan et al., 2005) as the source of the synthetic scenarios as it contains spontaneous meeting data which fits well with our target scenarios. AMI is an English multi-party meeting dataset with time-aligned transcripts and multiple microphone setups. As we mainly focus on a single main speaker and one bystander, the individual headset microphone (IHM) recordings were used, which provide per-speaker close-talk audio. We selected segments whose transcripts are complete and whose audio quality passes a basic SNR check, and retain the original speaker IDs and transcripts for role assignment.
Generating the Synthetic Scenarios: We first find audio segments of 2-3 minutes long where the speaker is speaking more than 70% of the time based on the rich transcription provided in AMI. These segments are used as the main speaker. We then find segments of 20-50 seconds with dense information content [^3] Segments containing substantive speech rather than backchanneling (e.g., “yes”, “uh-huh”). as the bystander audio segments. We attenuate bystander audio by -10dB relative to the main speaker, reflecting practical scenarios where bystanders are typically further from the microphone, and mix the bystander audio into the main segment at a random point. Meeting room reverberation is added during the mixing process.
@@
Given the scripts or transcripts of each audio file, we used GPT-4o to generate two types of annotations: (1) ten MCQs, five about the main speaker and five about the bystander, and (2) a content-conditioned natural language description of the main speaker. Each MCQ includes one correct answer, three distractors, and an additional “I don’t know”(IDK) option [^4] Note that we paraphrase the “I don’t know” into many different forms such as “I have no information” or “I cannot answer your question” to allow a more versatile test.. This option is essential, as a model with access only to the main speaker’s voice should select this option when asked about bystander content. Answer choices are randomly shuffled to reduce positional bias. The prompts used for annotation are provided in Appendix A.4.
@@ 4 Bystander Privacy Fine-Tuning
As a mitigation method, bystander privacy fine-tuning (BPFT) is proposed in this paper with a training pipeline specifically targeting the bystander privacy protection aspects. The goal of BPFT is to ensure that the model refuses to answer bystander-related questions when instructed to do so, while not losing the ability to comprehend speech content in a multi-speaker environment.
Specifically, following the same data creation pipeline for the synthetic audio partition (§3.2), we construct a larger training set containing 3,768 audio mixtures with 75k questions, including both MCQs and open-ended questions, with 37.5k related to the main speaker and 37.5k to the bystander. Each question in the training set has a pair of instructions, where one is to answer questions in general and another is to refuse if the question is about the bystander. As a result, this process will not only encourage the model to learn to distinguish and protect bystander privacy, but also enhance its performance in multi-speaker scenarios in general.
We selected Qwen-2.5-omni 7B (Xu et al., 2025) and Step-Audio-2-mini Boyong Wu et al. (2025) as two example open-source models to show the effectiveness of BPFT in bystander privacy protection. For training, we only fine-tune the LLM backbone with low-rank adaptation (LoRA) Hu et al. (2022) with rank 32, and freeze all other parts of the models.
@@ 5.1 Models
We thoroughly tested SH-Bench with a range of systems, including a pipeline system, open-source LLMs and mainstream proprietary models.
Pipeline System: A pipeline system comprises of a speech separation module to separate the main speaker out from other background audio, a speech recognition module to transcribe the speech into text, and an LLM to perform question-answering based on the transcriptions. Specifically we use SepFormer Subakan et al. (2021) to perform source separation, Whisper-Large-v3 Radford et al. (2023) for speech recognition and GPT-4o as the LLM to answer questions. Privacy-related instructions are provided through GPT-4o.
Open-source audio LLMs: We investigate popular LLMs with multi-speaker audio understanding abilities, including Qwen-2.5-Omni 3B and 7B Xu et al. (2025), Llama-Omni2-14B Fang et al. (2025), Step-Audio-2-mini Boyong Wu et al. (2025) and Kimi-Audio-7B-Instruct KimiTeam (2025). As a screening process, models were instructed to count how many speakers are present in the audio, and the above models were selected that can give the correct number most of the time. We closely followed the instructions provided on the official repositories, including the same system prompt and suggested decoding configurations for each model. Inference code is provided at https://github.com/Elocinacademia/SelectiveHearing-Bench.
Proprietary audio LLMs: We selected Gemini 2.5 Pro Comanici Gheorghe et al. (2025) and GPT-4o-audio-preview Hurst, Aaron et al. (2024) as two powerful proprietary models with audio perception abilities.